Bill weighs 201 lbs and has a maximum aerobic power of 40.7 ml/kg/min. What exercise intensity (power output in watts) on a cycle ergometer will enable him to maintain a target training intensity of 7.7 METs?
Answer #1

The equation for leg ergometry is given by

VO2 = 1.8*(Work/body mass)+7

Here, the target training intensity = 7.7 MET = 7.7*3.5 = 26.95 ml/kg/min

Body mass = 201 lbs = 91.172 kg

So, the work rate in watts is

26.95 = 1.8*(Work/91.172)+7

Work = (26.95-7)*91.172/1.8 = 1010.489 Watts